OpenStudy (anonymous):

verify (tanx+1)^2 =(1+sin2x)sec^2x

OpenStudy (anonymous):

(tanx+1)^2 = tan^2x + 2tanx + 1 = sin^2x + 2sinx + 1 ----- ----- cos^2x cos x = sin^2x + 2sinxcosx + cos^2x ------------------------- cos^2 x now use the identities sin^2x + cos^2x = 1 and 1/cos^2x = sec^2x can you continue?

OpenStudy (anonymous):

oh! also 2sinxcosx = sin2x

OpenStudy (anonymous):

Thank you so much for replying! I am still kind of confused about how you got to your third step. Would you mind explaining it to me?

OpenStudy (anonymous):

oh ok - we are converting to one fraction the LCM of cos^2x and cos x = cos^2x so we multiply each fraction by this LCM eg sin^2x cos^2x = sin^2x ------ * cos^2^x

OpenStudy (anonymous):

Thank you! So from the step you left off at sin^2x+2sinxcosx+cos^2x / cos^2x. s=I used the pythagorean identities sin^2x+cos^2x to equal 1. so know I have 1+2sinxcosx / cos^2x. If this correct? If so I'm not sure where to go from here

OpenStudy (anonymous):